# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, skipabove=10pt, skipbelow=10pt

### v1

@@ I Introduction
Co-optimizing safety and performance is a critical challenge in the design of controllers for autonomous systems, especially in safety-critical domains such as autonomous vehicles, UAVs, and robotics. In such settings, controllers must ensure that performance objectives (such as efficiency, speed, or agility) are met while guaranteeing that safety constraints are never violated. Achieving this balance can be viewed as a constrained optimal control problem, where the goal is to satisfy safety constraints throughout the trajectory while optimizing performance objectives.
@@
Several methods have been proposed to address this co-optimization challenge. Dynamic programming (DP) offers a rigorous solution to constrained optimization [1, 2]. Still, it is computationally intractable for many real-time applications due to the “curse of dimensionality”, often associated with DP-based solutions. On the other hand, MPC is more feasible for real-time applications, leveraging its ability to generate optimized control sequences over a receding horizon [3, 4]. Among these, Model Predictive Path Integral (MPPI) control is a sampling-based MPC method that has gained significant popularity due to its flexibility in handling complex dynamics, uncertainties, and non-linear systems [5, 6]. MPPI leverages stochastic sampling to optimize control trajectories, offering a scalable and efficient way to generate control actions. However, ensuring safety within MPPI has proven challenging, as the basic framework does not inherently account for hard safety constraints. Consequently, various extensions have been proposed to incorporate safety, each solving the co-optimization problem differently.
@@
Classical MPPI-based approaches encourage safety constraints by penalizing unsafe sampled trajectories in the cost function, such that they are effectively ignored in the sample aggregation process. However, this approach cannot guarantee safety and its satisfaction depends on the penalty function. Another drawback is that computation is wasted on ignored samples, reducing the effective number of samples used for optimization. To enforce safety, approaches such as safety filtering and barrier functions have emerged, where constraints are enforced as a post-optimization correction [7, 8]. Such approaches solve the MPPI problem first and apply safety filtering afterward, or combine these two approaches [9, 10, 11].However, these approaches tend to ignore the long-term effect of safety action on the performance. Moreover, the overall performance of the system is still limited by the underlying MPPI algorithm. Other procedures use probabilistic methods to address the safety-performance co-optimization, such as modifying the sampling distribution to favor safer trajectories [12], using stochastic safety certificates to steer away from unsafe regions [13], or improving safety indirectly by enhancing adaptability and robustness to uncertainty [14, 15].
@@
In this work, we introduce DualGuard-MPPI, a novel MPPI algorithm that incorporates safety constraints directly into the MPPI algorithm using Hamilton-Jacobi (HJ) reachability. Our key idea is to incorporate two safety filtering stages in the MPPI algorithm: first, we perform safety filtering along the sampled control perturbations, leading to provably safe “hallucinations” of the system trajectories, which are then used to optimize the performance objective. Second, we apply safety filtering on the selected control sequence before it is applied to the system to filter potential multi-modal sequences that may compromise safety when combined. The proposed approach imposes safety constraints throughout the trajectory optimization and control application stages, ensuring that the system operates within safe bounds at all times, without requiring any safety penalty tuning. Moreover, since all generated samples are provably safe, they all contribute to performance optimization, effectively reducing the MPPI sampling variance and leading to a better performance, as compared to vanilla MPPI algorithm.
In summary, the key contributions of this work are:
@@
  • An MPPI framework that co-optimizes safety and performance while ensuring safety at all times;
@@ III-A Model Predictive Path Integral (MPPI)
MPPI [5, 6] aims to solve a classical stochastic optimal control problem using a zeroth-order sampling-based MPC scheme. Specifically, at each state, MPPI samples K random control sequences around a nominal control sequence for a time horizon $\hat{T}\leq t_{f}$ discretized over $H$ evenly spaced steps. The control input for the kth sequence at jth time step is given by: $u^{k}(x,j)=u(x,j)+\delta_{j}^{k}$, where $u(x,j)$ is a nominal control sequence and $\delta_{j}^{k}$ a randomly sampled control perturbation. Next, the state trajectory $\xi_{j}^{k}$ and the cost to go $S(\xi_{j}^{k})$ corresponding to each control sequence is computed. It is shown that the optimal control sequence can be approximated by the following update law [9]:
@@
Dotted lines in Fig. 1 represent possible rollouts of the system following a randomly perturbed control sequence starting from the current state. We refer to these as the hallucinations of the system for a given state.
Figure 1: Visualization of hallucination step in the MPPI algorithm.
The hallucinations that fail to avoid obstacles achieve a higher cost (shown in red). Thus, the update law in (3) prioritizes control perturbations that will keep the car safe (shown in a blue-to-purple gradient, with blue indicating lower costs). However, in general, it could be possible that all samples lead to safety violations. Moreover, if the number of safe samples is small, it could lead to a high variance in the performance of MPPI, necessitating the need to account for safety constraints rigorously.
@@ III-B Hamilton-Jacobi Reachability
In safety-ensuring applications, we want to guarantee that the system does not enter $\mathcal{F}$ at any time; for this reason, we use a time-converged BRT, as the set of unsafe states often stops growing after some amount of time. Consequently, we can use the converged value function $V(\mathbf{x})$, which, when used to synthesize safety controllers, gives an expression identical to (10) without its time dependency.
@@ III-C Least Restrictive Filtering
We can use the HJI-VI in (6) to calculate the value function and associated BRT, over which we can implement a LRF to guarantee the system’s safety. In this case considering the dynamics in (4) and the environment definition, we compute the value function using the LevelSetToolbox[21].
In Fig. 2, the purple trajectory represents the system’s response to a perturbed control sequence, reaching a critical state at the boundary of the safe set (yellow dot). The boundary of the safe set at this state is illustrated with a dotted line. The red trajectory shows the continuation of the system using the perturbed controls, which leads to the failure set. By contrast, the blue trajectory demonstrates how the unsafe execution can be made safe. This is achieved by applying a LRF step to each control action in the sequence before execution.
@@ IV Dualguard MPPI
As discussed in Sec. III-A, MPPI solves the optimal control problem in (2) using a sampling-based method. While MPPI may encourage the satisfaction of safety constraints in (2) via penalizing safety violations in the cost function, ensuring safety constraints remains challenging. Moreover, by construction, the high-cost safety breaching sequences are mostly ignored during the optimal control sequence computation, wasting computational resources that could have been used to refine the system performance further. While the safety filtering mechanism discussed in III-C can provide a safety layer after MPPI to enforce the safety constraint, this approach is myopic in nature and might lead to performance impairment in favor of safety.
To overcome these challenges, we propose DualGuard MPPI, a two-layered safety filtering approach. First, the safety filtering is incorporated during the sampling process itself, where a least restrictive filter is applied along the sampled control sequence rollouts using a pre-computed safety value function. This ensures that all hallucinations satisfy the safety constraint and can contribute to performance optimization. To ensure that the resultant optimal control sequence also satisfies the safety constraint, we filter the output optimal control sequence by a safety filter as well. In addition to ensuring the safety constraints, the proposed framework leads to an increased sample efficiency as the safe hallucinations keep all sampled trajectories safe and relevant to the performance objective, thereby avoiding ”sample wastage” due to safety constraint violation. The proposed algorithm is presented in Alg. 1. Details on the proposed filtering stages are discussed next.
IV-A Generating Safe hallucinations
As in classical MPPI, we consider $K$ sequences of random perturbations $\delta_{j}^{k}$, that modify a nominal control sequence $u_{j}$. The perturbed sequences are applied to the dynamic model of the system while being filtered at each time step along the horizon, using the LRF in eq. (11), resulting in hallucinated trajectories that are guaranteed to maintain safety. The cost-to-go $S^{k}$ for the filtered trajectory is calculated over the safety-filtered control perturbation sequence $\Delta_{j}^{k}$.
The nominal control sequence $u_{j}$, the filtered control perturbations $\Delta_{j}^{k}$ and the cost over the safe hallucinated trajectories $S^{k}$ are used to calculate the optimal control sequence $u^{*}_{j}$ using the update rule defined in (3). As we weigh over controls that only produce safe trajectories, all K sequences contribute to the performance optimization, reducing the variance of MPPI algorithm and leading to better performance.
@@ IV-B Output Least restrictive filtering
One case that encourages this last filtering stage corresponds to scenarios where the safe hallucinations present multimodality. We can imagine a vehicle trying to swerve around an obstacle by turning right or left; even if both swerving maneuvers result in safe behavior, a weighting over hallucinations split between these two modalities could result in maintaining a straight trajectory, causing a collision. Such issues are resolved by this additional filtering stage which picks one of the safe modes.
@@
Considering our running example we observe the proposed steps of DualGuard MPPI in Fig. 3. First, we visualize the unmodified hallucination step in the left panel. Next, using the same control perturbations, we show how the safe hallucination step renders all the samples safe. Hallucinated trajectories are presented in a blue-to-red scale, corresponding to the associated low-to-high costs.
The right panel of Fig. 3 shows how the safe hallucinations might present themselves in a multimodal fashion for some obstacle configurations; this indicates a possible failure mode where the control sequence given by the update rule in (3) could drive the vehicle directly into the obstacles, leading to safety violations. The proposed output filtering stage safeguards against this possibility.
Figure 3: Unmodified hallucinations (Left). Safe hallucinations (Center). Possible multimodality on safe hallucinations encourages the use of output least restrictive filtering (Right).
@@ V Simulation Studies
We now evaluate the performance of our proposed method across several simulation studies and on a real hardware testbed. For this, we compare the proposed approach against five baseline methods. Each baseline builds on the MPPI control approach but incorporates distinct modifications to enhance safety.
@@
BRT Penalty (BRT penalty): Advanced MPPI methods penalize hallucinated trajectories that enter a safe zone defined as augmented obstacles, Discrete Barrier States (DBaS), or Control Barrier Functions (CBF) [25, 26, 27, 28]. To contrast against this technique, we implement a baseline that penalizes hallucinated trajectories that enter the BRT of the obstacles. This approach is more preemptive, as it penalizes samples that have entered states from which such a collision is unavoidable but may have not yet collided with obstacles within the sampling horizon.
Obstacle Penalty with Output LRF (Obs. pen. + LRF): Extension of the Obstacle Penalty baseline by adding a least restrictive filtering step. This corresponds to a naive safety filtering approach where the least restrictive filter is used to correct any control that would result in a safety violation just before it is applied.
@@
Shield Model Predictive Path Integral (Shield MPPI): This method leverages the approach described in Shield-MPPI[10], where a cost related to the CBF condition is considered during hallucinations, followed by an approximate CBF-based repair step over the resulting updated control to promote safety. To ensure a fair comparison, we will use the HJ value function as a barrier function so that all methods work on the maximal safe set [29].
Evaluation Metrics: To quantitatively assess the performance of each method, we define the following evaluation metrics. These metrics will be measured over batches of simulation runs for each method, providing insights into safety, efficiency, and computational demands.
@@
  • RelCost: Average normalized cost of the trajectory, using the proposed method as the baseline. We only consider successful executions across all methods for RelCost computation. This allows for a fair comparison and ensures that the RelCost is not affected by the high failure penalties for some of the baselines.
@@ V-A Safe Planar Navigation
We assess the running-example safe planar navigation with dynamics defined in (4). In each episode, the system must navigate from an initial state $\mathbf{x}_{0}$ to a goal region of radius 0.1 m in the $xy$ plane centered at the goal state $\mathbf{x}_{g}$, within a time horizon of $T=20\text{ s}$. The optimal control problem’s cost function is defined as the squared distance from the goal, a control effort penalty, and a safety penalty $P(\mathbf{x})$:
@@
with $Q=\text{diag}([1,1,0])$, $m=0.2$, and $P=10000$ if the state is inside the obstacles or the BRT (corresponding to that controller’s cost penalty type) and $K=0$ otherwise.
We test each controller in the same set of 100 episodes, where each episode is defined by a particular pair $(\mathbf{x}_{0},\mathbf{x}_{g})$. These initial-goal state pairs were selected randomly but constrained to be outside the BRT, near the environment boundary, and separated by at least 5 m for diversity. The results are summarized in Table I. In addition to the metrics previously defined, we also compute the Success rate, defined as the percentage of episodes that reach the goal region safely (i.e., without collisions) within the time horizon, and Timeout, defined as the percentage of episodes that fail to reach the target within the given time horizon.
@@
TABLE I: Safe Planar Navigation results over 100 episodes.
@@
1000 Obs. penalty 49 1 50 1.34 ± 1.19
BRT penalty 72 21 7 1.89 ± 2.11
Obs. pen. + LRF 80 20 0 1.35 ± 1.20
BRT pen. + LRF 74 26 0 1.90 ± 2.12
Shield MPPI 83 17 0 1.32 ± 0.87
DualGuard (Ours) 99 1 0 1.00
250 Obs. penalty 31 0 69 1.10 ± 0.24
BRT penalty 58 15 27 2.35 ± 3.40
Obs. pen. + LRF 81 19 0 1.27 ± 0.92
BRT pen. + LRF 69 31 0 1.86 ± 2.03
Shield MPPI 78 21 1 1.40 ± 0.86
DualGuard (Ours) 98 2 0 1.00
60 Obs. penalty 11 0 89 0.99 ± 0.03
BRT penalty 43 21 36 3.16 ± 4.17
Obs. pen. + LRF 70 30 0 0.99 ± 0.03
BRT pen. + LRF 55 45 0 3.16 ± 4.17
Shield MPPI 64 36 0 1.08 ± 0.14
DualGuard (Ours) 96 4 0 1.00
First examining the results for $K=1000$, we see that our proposed method achieves the lowest failure rate and highest success rate. While all methods with an output Least Restrictive Filter avoid collisions as expected, our method also times-out in fewer episodes, indicating a better performance. The safety-performance co-optimization capabilities of the proposed method are also evident from its lowest RelCost.
To understand the different controllers’ behavior, consider the trajectories shown in Figure 4. Here, the obstacle penalty method fails early by entering a region in the northeast corner which is impossible to escape and hence avoided by all BRT-aware methods. The other baseline methods execute a reasonable trajectory to the goal, but the proposed method manages to find a lower-cost trajectory by taking a shortcut south after rounding the first obstacle. This is because the baseline methods are still fundamentally limited by the exploration capabilities of MPPI – if the underlying samples are unsafe, they contribute little to the performance optimization. On the other hand, the proposed method generates safe hallucinations, allowing it to explore more relevant parts of the environment with the same number of samples, as previously visualized in Fig. 3, leading to a better performance optimization.
A consequence of the superior exploration capabilities of the proposed method is that it can achieve similar success rates with far fewer samples. We summarize the results for different methods as we vary the number of samples in Table I. We note that while all methods degrade with fewer samples as expected, DualGuard remains relatively consistent. With only 60 samples, it succeeds in 96% of the episodes, with the closest contender not going over 70%. This could enable the deployment of the proposed method on resource-constrained systems, which MPPI methods typically struggle with (as also indicated by the poor performance of other MPPI baselines with fewer samples).

### v2

@@ Abstract
Index Terms:
Robot Safety, Collision Avoidance, Motion and Path Planning.
@@ I Introduction
Co-optimizing safety and performance is a critical challenge in the design of controllers for autonomous systems, especially in safety-critical domains such as autonomous vehicles, UAVs, and robotics. In such settings, controllers must ensure that performance objectives are met while guaranteeing that safety constraints are respected. This can be framed as a constrained optimal control problem, where the goal is to satisfy safety constraints throughout the trajectory while optimizing performance objectives.
@@
Several methods have been proposed to address this co-optimization challenge. Dynamic programming (DP) offers a rigorous solution to constrained optimization [1, 2]. Still, it is computationally intractable for many real-time applications due to the “curse of dimensionality”. On the other hand, MPC is more feasible for real-time applications, leveraging its ability to generate optimized control sequences over a receding horizon [3, 4]. Among these, Model Predictive Path Integral (MPPI) control is a sampling-based MPC method that has gained popularity due to its flexibility in handling complex dynamics, uncertainties, and non-linear systems [5]. MPPI leverages stochastic sampling to optimize control trajectories, offering a scalable and efficient way to generate control actions. However, ensuring safety within MPPI has proven challenging, as the basic framework does not account for hard safety constraints. Consequently, various extensions have been proposed to incorporate safety, each solving the co-optimization problem differently.
@@
Classical MPPI-based approaches encourage safety by penalizing unsafe sampled trajectories, which are then mostly ignored in the sample aggregation process. However, this approach cannot guarantee safety and its satisfaction depends on the penalty function. Another drawback is that computation is wasted on ignored samples, reducing the effective number of samples used for optimization. To enforce safety, approaches such as safety filtering and barrier functions have emerged, where constraints are enforced as a post-optimization correction [6, 7]. Such approaches allow to solve the MPPI problem first and apply safety filtering afterward, or combine these two approaches [8, 9, 10]. However, these approaches tend to ignore the long-term effect of safety actions on performance. Other procedures use probabilistic methods to address the safety-performance co-optimization, such as modifying the sampling distribution to favor safer trajectories [11], using stochastic safety certificates to steer away from unsafe regions [12], or improving safety indirectly by enhancing adaptability and robustness to uncertainty [13, 14].
Beyond MPPI literature, control barrier function methods, allow for the co-optimization of safety and performance by solving constrained optimization problems [7, 15]. However, these approaches tend to be myopic, optimizing safety and performance locally without capturing long-term effects, and often rely on Lyapunov-based objectives that limit the expressiveness of performance goals. Filter-aware motion planning has also been explored to mitigate this performance degradation caused by safety interventions, where instead of treating safety reactively, these approaches explicitly account for future interventions. For instance, shielding-aware planning proactively balances performance with the risk of emergency maneuvers triggered by low-probability agent behaviors [16, 17]. Similarly, in reinforcement learning settings, integrating safety filters during training rather than solely during execution has been shown to improve performance and reduce undesired safety interventions[18, 19].
@@
In this work, we introduce DualGuard-MPPI, a novel MPPI algorithm that incorporates safety constraints directly into the MPPI algorithm using Hamilton-Jacobi (HJ) reachability. Our key idea is to incorporate two safety filtering stages in the MPPI algorithm: first, we perform safety filtering along the sampled control perturbations, leading to provably safe rollouts of the system trajectories, which are then used to optimize the performance objective. Second, we apply safety filtering on the selected control sequence before it is applied to the system to prevent multimodal combinations that may compromise safety. In this context, multimodality refers to distinct groups of safe trajectories, where different evasive maneuvers—such as swerving left or right to avoid an obstacle—can both be viable but should not be arbitrarily mixed.
The proposed approach imposes safety constraints throughout the trajectory optimization and control application stages, ensuring that the system operates within safe bounds at all times, without requiring any safety penalty tuning. Moreover, since all generated samples are provably safe, they all contribute to performance optimization, effectively reducing the MPPI sampling variance and leading to a better performance, as compared to vanilla MPPI algorithm. In summary, the key contributions of this work are:
@@
  • An MPPI framework that co-optimizes safety and performance while ensuring safety at all times.
@@ III-A Model Predictive Path Integral (MPPI)
Original MPPI [5] aims to solve a stochastic optimal control problem using a zeroth-order sampling-based MPC scheme for undisturbed deterministic general nonlinear systems. Specifically, at each state, MPPI samples K random control sequences around a nominal control sequence for a time horizon $\hat{T}\leq t_{f}$ discretized over $H$ evenly spaced steps. The control input for the kth sequence at jth time step is given by: $u^{k}(x,j)=u(x,j)+\delta_{j}^{k}$, where $u(x,j)$ is a nominal control sequence and $\delta_{j}^{k}$ a randomly sampled control perturbation. Next, the state trajectory $\xi_{j}^{k}$ and the cost to go $S(\xi_{j}^{k})$ corresponding to each control sequence is computed. It is shown that the optimal control sequence can be approximated by the following update law [8]:
@@
Dotted lines in Fig. -1 represent possible rollouts of the system following a randomly perturbed control sequence starting from the current state.
Figure -1: Visualization of sampling step in the MPPI algorithm.
The rollouts that fail to avoid obstacles achieve a higher cost (shown in red). Thus, the update law in (3) prioritizes control perturbations that will keep the car safe (shown in a blue-to-purple gradient, with blue indicating lower costs). However, it could be possible that all samples lead to safety violations. Moreover, if the number of safe samples is small, it could lead to a high variance in the performance of MPPI, necessitating the need to account for safety constraints rigorously.
@@ III-B Hamilton-Jacobi Reachability
In safety-ensuring applications, we want to guarantee that the system does not enter $\mathcal{F}$ at any time. To achieve this, we use a time-converged BRT, as the set of unsafe states typically stops expanding after some time. Since our objective is safety, we will rely exclusively on the converged value function $V(\mathbf{x})$ throughout this paper. This allows us to synthesize safety controllers using an expression identical to (10) but without time dependency.
@@ III-C Least Restrictive Filtering
We can use the HJI-VI in (6) to calculate the value function and associated BRT, over which we can implement a LRF to guarantee the system’s safety. In this case considering the dynamics in (4) and the environment definition, we numerically compute the value function using the LevelSetToolbox [25].
In Fig. III-C, the purple trajectory represents the system’s response to a perturbed control sequence, reaching a critical state at the boundary of the safe set (yellow dot). The boundary of the safe set at this state is illustrated with a dotted line. The red trajectory shows the continuation of the system using the perturbed controls, which leads to the failure set. By contrast, the blue trajectory demonstrates how the unsafe execution can be made safe. This is achieved by applying a LRF step to each control action in the sequence before execution.
@@ IV Dualguard MPPI
As discussed in Sec. III-A, MPPI addresses the optimal control problem in (2) using a sampling-based method. While MPPI may encourage the satisfaction of safety constraints in (2) via penalizing safety violations in the cost function, ensuring safety constraints remains challenging. Moreover, by construction, the high-cost safety breaching sequences are mostly ignored during the optimal control sequence computation, wasting computational resources that could have been used to refine the system performance further. While the safety filtering mechanism discussed in III-C can provide a safety layer after MPPI to enforce the safety constraint, this approach is myopic in nature and might lead to performance impairment in favor of safety.
To overcome these challenges, we propose DualGuard MPPI, a two-layered safety filtering approach. Offline, we use HJ reachability to compute a safety value function considering the disturbed system dynamics. By incorporating disturbances at this stage, we ensure safety against worst-case scenarios. In the first step of our algorithm, safety filtering is incorporated directly into the original MPPI framework, where control sequence rollouts sampled using the undisturbed dynamics are rendered safe by a LRF step. Disturbances are not included during MPPI propagation to streamline the sampling process, since safety constraints are already handled by the LRF. This ensures that all rollouts remain within the safe set and can fully contribute to performance optimization. To ensure that the resultant optimal control sequence also satisfies the safety constraint, we filter the output optimal control sequence by a safety filter as well. In addition to ensuring the safety constraints, the proposed framework leads to an increased sample efficiency as the safe rollouts keep all sampled trajectories safe and relevant to the performance objective, thereby avoiding “sample wastage” due to safety constraint violation. The proposed algorithm is presented in Alg. 1. Details on the proposed filtering stages are discussed next.
IV-A Generating Safe Rollouts
As in classical MPPI, we consider $K$ sequences of random perturbations $\delta_{j}^{k}$, that modify a nominal control sequence $u_{j}$. The perturbed sequences are applied to the deterministic dynamic model of the system while being filtered at each time step along the horizon, using the LRF in eq. (11), resulting in rollouts that are guaranteed to maintain safety. The cost-to-go $S^{k}$ for the filtered trajectory is calculated over the safety-filtered control perturbation sequence $\Delta_{j}^{k}$.
The nominal control sequence $u_{j}$, the filtered control perturbations $\Delta_{j}^{k}$ and the cost over the safe rollouts $S^{k}$ are used to calculate the optimal control sequence $u^{*}_{j}$ using the update rule defined in (3). As we weigh over controls that only produce safe trajectories, all K sequences contribute to the performance optimization, reducing the variance of MPPI algorithm and leading to better performance.
@@ IV-B Output Least restrictive filtering
One motivation for this last filtering stage is scenarios where the safe rollouts present multimodality. We can imagine a vehicle trying to swerve around an obstacle by turning right or left; even if both maneuvers result in safe behaviors, a weighting over rollouts split between these modalities could result in maintaining a straight trajectory, causing a collision. Such issues are resolved by this additional filtering stage which picks one of the safe modes.
@@
Considering our running example we observe the proposed steps of DualGuard MPPI in Fig. 1. First, we visualize the unmodified rollout step in the left panel. Next, using the same control perturbations, we show how the safe rollout step renders all the samples safe. Sampled trajectories are presented in a blue-to-red scale, corresponding to the associated low-to-high costs.
The right panel of Fig. 1 shows how the safe rollouts might present themselves in a multimodal fashion for some obstacle configurations; this indicates a possible failure mode where the control sequence given by the update rule in (3) could drive the vehicle directly into the obstacles, leading to safety violations. The proposed output filtering stage safeguards against this possibility.
Figure 1: Unmodified rollouts (Left). Safe rollouts (Center). Possible multimodality on safe rollouts encourages the use of output least restrictive filtering (Right).
@@ V Simulation Studies
We evaluate the performance of our proposed method on simulation studies and a real hardware testbed. For this, we compare the proposed approach against five baseline methods. Each baseline builds on the MPPI control approach but incorporates distinct modifications to enhance safety.
@@
BRT Penalty (BRT penalty): Advanced MPPI methods penalize sampled trajectories that enter a safe zone defined as augmented obstacles, Discrete Barrier States (DBaS), or Control Barrier Functions (CBF) [29, 30, 31]. To represent this technique, we implement a baseline that penalizes trajectories that enter the obstacle set BRT. This preemptive approach penalizes trajectories that have entered states from which collision is unavoidable, even if they have not yet entered the obstacle set within the sampling horizon.
Obstacle Penalty with Output LRF (Obs. pen. + LRF): Extension of the Obstacle Penalty baseline by adding a LRF step. This corresponds to a naive safety filtering approach where the LRF is used to correct any control that would result in a safety violation just before it is applied.
@@
Shield Model Predictive Path Integral (Shield MPPI): We implement the approach described in [9], where a cost related to the CBF condition is considered for rollouts, followed by an approximate CBF-based repair step over the updated control to promote safety. To ensure a fair comparison, we use the HJ value function as a barrier function so that all methods work on the maximal safe set [32].
Evaluation Metrics: To quantitatively assess the performance of each method, we define the following evaluation metrics. These metrics will be measured over batches of simulation runs for each method.
@@
  • RelCost: Average cost of the trajectories and standard error, normalized to our method’s mean cost. We only consider common safe trajectories between the compared method and ours. This allows for a fair comparison and ensures that the RelCost is not affected by the high failure penalties for some of the baselines.
@@ V-A Safe Planar Navigation
We assess the running-example safe planar navigation with dynamics defined in (4). In each episode, the system must navigate from an initial state $\mathbf{x}_{0}$ to a goal region of radius 0.1 m in the $xy$ plane centered at the goal state $\mathbf{x}_{g}$, within a time horizon of $T=20\text{ s}$. The cost function is defined as the squared distance from the goal, a control effort penalty, and a method dependent safety penalty $P(\mathbf{x})$:
@@
We test each controller in the same set of 100 episodes defined by pairs $(\mathbf{x}_{0},\mathbf{x}_{g})$. These initial-goal state pairs were selected randomly but constrained to be outside the BRT, near the environment boundary, and separated by at least 5 m for diversity. The results are summarized in Table I. In addition to the metrics previously defined, we also compute the Success rate, defined as the percentage of episodes that reach the goal region safely within the time horizon, and Timeout, defined as the percentage of episodes that fail to reach the target within the given time horizon.
@@
TABLE I: Safe Planar Navigation results over 100 episodes. (*) Indicates statistically significant differences ($p<0.05$) compared to our method.
@@
1000 Obs. penalty 49 1 50 1.19 ± 0.14*
BRT penalty 72 21 7 2.05 ± 0.17*
Obs. pen. + LRF 80 20 0 1.44 ± 0.10*
BRT pen. + LRF 74 26 0 2.61 ± 0.29*
Shield MPPI 83 17 0 1.70 ± 0.14*
DualGuard (Ours) 99 1 0 1.00 ± 0.07
250 Obs. penalty 31 0 69 0.99 ± 0.10
BRT penalty 58 15 27 2.28 ± 0.23*
Obs. pen. + LRF 81 19 0 1.80 ± 0.14*
BRT pen. + LRF 69 31 0 5.46 ± 0.59*
Shield MPPI 78 21 1 1.85 ± 0.14*
DualGuard (Ours) 98 2 0 1.00 ± 0.08
60 Obs. penalty 11 0 89 1.07 ± 0.16
BRT penalty 43 21 36 3.70 ± 0.49*
Obs. pen. + LRF 70 30 0 2.45 ± 0.18*
BRT pen. + LRF 55 45 0 6.48 ± 0.60*
Shield MPPI 64 36 0 2.82 ± 0.20*
DualGuard (Ours) 96 4 0 1.00 ± 0.07
Examining the results for $K=1000$, we see that DualGuard achieves zero failure rate and highest success rate. While all methods with an output LRF avoid collisions as expected, our method also times-out in fewer episodes, indicating a better performance. The safety-performance co-optimization capabilities of the proposed method are also evident from its lowest RelCost across the safe methods, this difference was further validated to be statistically significant through paired t-tests against DualGuard-MPPI.
To understand the different controllers’ behavior, consider the trajectories shown in Figure 2. Here, the obstacle penalty method fails early by entering a region in the northeast corner which is impossible to escape and hence avoided by all BRT-aware methods. The other baseline methods execute a reasonable trajectory to the goal, but the proposed method manages to find a lower-cost trajectory by taking a shortcut after rounding the first obstacle. This is because the baseline methods are still fundamentally limited by the exploration capabilities of MPPI – if the underlying samples are unsafe, they contribute little to the performance optimization. On the other hand, the proposed method generates safe rollouts, allowing it to explore more relevant parts of the environment with the same number of samples, as previously visualized in Fig. 1, leading to a better performance optimization.
A consequence of the superior exploration capabilities of the proposed method is that it can achieve similar success rates with far fewer samples. We summarize the results for different methods as we vary the number of samples in Table I. We note that while all methods degrade with fewer samples as expected, DualGuard remains relatively consistent. With only 60 samples, it succeeds in 96% of the episodes, with the closest contender not going over 70%. This could enable its deployment on resource-constrained systems, which MPPI methods typically struggle with.